### v2

# Let Guidelines Guide You: A Prescriptive Guideline-Centered Data Annotation Methodology

Abstract
We introduce the Guideline-Centered annotation process, a novel data annotation methodology focused on reporting the annotation guidelines associated with each data sample. We identify three main limitations of the standard prescriptive annotation process and describe how the Guideline-Centered methodology overcomes them by reducing the loss of information in the annotation process and ensuring adherence to guidelines. Additionally, we discuss how the Guideline-Centered enables the reuse of annotated data across multiple tasks at the cost of a single human-annotation process.
Trigger Warning: This paper contains examples of offensive and explicit content.
1 Introduction
In recent years, several works have highlighted the importance of clear guidelines for producing high-quality data annotations. In particular, researchers have framed distinct annotation processes into annotation paradigms, each with its own set of characteristics, limitations, and requirements. The most notable contributions in this field are the prescriptive and descriptive annotation paradigms, originally designed to address subjective tasks in NLP Rottger et al. (2022), such as hate speech detection, subjectivity detection, and sentiment analysis. The prescriptive annotation paradigm enforces a shared belief of the task conveyed into a unique set of annotation guidelines on which annotators are instructed. In contrast, the descriptive annotation paradigm favors annotators’ perspectives and beliefs, which are allowed to be reflected in the task. Another notable emergent annotation paradigm is the perspectivist one Cabitza et al. (2023), where multiple annotation ground truths for the same data sample are allowed while sharing a unique set of annotation guidelines. All these paradigms share the underlying motivation of addressing annotation issues, which lead to low-quality data and are largely evident in subjective tasks like hate speech detection Plank et al. (2014); Sap et al. (2022), and subjectivity detection Riloff and Wiebe (2003); Savinova and Moscoso Del Prado (2023).
Term (symbol) Definition
Annotation Process of mapping data samples to a class set $\mathcal{C}$.
Class set ($\mathcal{C}$) Set of classes (i.e., symbols) by which to annotate.
Guideline set ($\mathcal{G}$) Set of textual guideline definitions that describe a task.
Guideline ($g$) Textual definition describing a particular aspect of a task.
Data sample ($x$) Textual sample to be annotated.
Guideline subset ($\mathcal{G}_{x}$) Guideline subset associated with sample $x$.
Class subset ($\mathcal{C}_{x}$) Class subset associated with sample $x$.
Standard decision-making function ($f$) Process of aligning a data sample $x$ to its class subset $\mathcal{C}_{x}$.
GC decision-making function ($h$) Process of aligning a data sample $x$ to its guideline subset $\mathcal{G}_{x}$.
GC class grounding function ($r$) Process of mapping a guideline definition $g\in\mathcal{G}_{x}$ to its class $c\in\mathcal{C}$.
Table 1: Glossary of the main concepts discussed in this article.
While each annotation paradigm aims at a specific intended use of the collected data and provides recommendations on how to coherently set the annotation process correspondingly, all instances of these annotation paradigms are centered on aligning data samples to a class set tailored to the specific task of interest. We argue that an annotation process relying on reporting class-set information for each data sample has several weaknesses that can notably affect the quality of collected data. We elaborate on our hypothesis by analyzing the prescriptive annotation paradigm, where convergence to the shared belief is an indicator of data quality and where a set of annotation guideline definitions is a mandatory requirement for data annotation Rottger et al. (2022).
We propose a shift in the annotation process, where annotators explicitly leverage the definitions delineated in validated annotation guidelines rather than relying on less informative class sets. In particular, we introduce a novel annotation methodology, denoted Guideline-Centered (GC), focused on designing an annotation process that adheres transparently to employed annotation guidelines. GC opens a wide spectrum of advantages that affect the way data is annotated and eventually used for model training.
The remainder of this paper is structured as follows. In Section 2, we discuss the standard annotation process under the prescriptive paradigm. Then, we introduce our methodology in Section 3. In Section 4 we frame our contribution with respect to the existing literature. Finally, we discuss some major implications of GC in Section 5, and report the concluding remarks in Section 6.
2 Background and Motivation
Figure 1: (a) Example of guideline set $\mathcal{G}$ and class set $\mathcal{C}$ provided for annotation concerning hate speech detection. (b) In the standard annotation process, the annotator Refer to caption aligns data sample $x$ to the class subset $\mathcal{C}_{x}$ after being instructed with the mapping between $\mathcal{G}$ and $\mathcal{C}$. (c) In our two-stage Guideline-Centered annotation process, first, the annotator Refer to caption aligns $x$ to the guideline subset $\mathcal{G}_{x}$ (blue box) and, then, the class subset $\mathcal{C}_{x}$ is retrieved from $\mathcal{G}_{x}$ using the class grounding function known apriori (green box). The example depicted is freely inspired by Kirk et al. (2023).
In a data annotation process performed according to the prescriptive paradigm Rottger et al. (2022), [^1] Inspired by Cabitza et al. (2023) we refer the reader to Table 1 to avoid potential ambiguities in the presented notation. a group of annotators labels data samples by following a set of detailed guidelines encoding the task. The labeling consists of mapping each data sample to one or more instances of the provided class set $\mathcal{C}$. The class set $\mathcal{C}$ is task-specific and may vary according to the chosen problem formulation. For instance, in hate speech detection Jahan and Oussalah (2023), annotators may distinguish between hate and non-hate examples ($\mathcal{C}=\{\texttt{hate},\,\texttt{non{-}hate}\}$) or identify a more fine-grained taxonomy ($\mathcal{C}=\{\texttt{racism},\,\texttt{misogyny},\,\dots\}$). Hereinafter, we refer to a data annotation process performed according to the prescriptive paradigm as standard annotation process or simply data annotation process. In the remainder of the section, we formalize the standard annotation process and elaborate on its limitations that motivate our work.
2.1 Standard Annotation Process Limitations
Formally, given a dataset $\mathcal{X}=\{x_{1},x_{2},\dots,x_{N}\}$ and a class set $\mathcal{C}=\{c_{1},c_{2},\dots,c_{K}\}$, the data annotation process consists of aligning a data sample $x\in\mathcal{X}$ to its specific class subset $\mathcal{C}_{x}\subseteq\mathcal{C}$ by relying on the guideline set $\mathcal{G}=\{g_{1},g_{2},\dots,g_{M}\}.$ [^2] In a multi-class classification setting, the sample-specific class subset contains only one class (i.e., $\mathcal{C}_{x}=\{c_{x}\}$). The annotators are provided with the mapping between the guideline set $\mathcal{G}$ and the class set $\mathcal{C}$ to ground their decision-making process. However, they are not instructed to report the subset of guidelines they rely on for annotating each data sample. For example, in Figure 1(b), the annotator implicitly uses the mapping between $\mathcal{G}$ and $\mathcal{C}$ to label $x$ as sexist but at the end of the annotation process we do not have access to the guidelines they relied on to make their decision. We see this loss of information as a major drawback of the standard annotation process, potentially undermining the overall annotation quality. In the remainder of the section, we discuss this aspect in detail, identifying three main critical issues.
Opaque Annotation Adherence to Guidelines.
A crucial step in the data annotation process is assessing the annotators’ qualification. Standard practice dictates that annotators should undergo a preliminary test to assess their eligibility. When prompted with some annotation tests, a good annotator should assign the correct class subset $\mathcal{C}_{x}$ to each data sample $x$ following the provided guidelines $\mathcal{G}$ as a measure of their understanding of the task. Verifying the adherence of the labeling process to the guideline set $\mathcal{G}$ is a desideratum in any annotation process as a measure of its validity. However, without having access to the guideline subset the annotator associates with each data sample, we cannot verify whether the annotator relied exclusively on the provided guidelines to annotate data or used their own internalized criteria.
Coarse-grained Annotation Evaluation.
In the prescriptive paradigm, a high inter-annotator agreement is an indicator of the quality of a dataset Rottger et al. (2022). Standard practice in data annotation dictates a preliminary stage aimed at an iterative refinement of the guideline set $\mathcal{G}$, which usually consists of one or multiple discussion phases on identified edge cases Caselli et al. (2020); Guzmán-Monteza (2023); Jikeli et al. (2023); Kirk et al. (2023); Antici et al. (2024). Each discussion should discourage annotator subjectivity in favor of a shared belief represented by the guideline set and highlight commonalities and differences among annotators’ points of view. However, the annotators’ discussion revolves around differences in the selected class subset $\mathcal{C}_{x}$, without an explicit reference to the guideline subset associated with sample $x$. This prevents the evaluation of annotators’ points of view under the lens of used guidelines and does not allow for a fine-grained evaluation of annotators’ agreement, which is limited to computing quality measures in terms of class subsets.
Standard Guideline-Centered
Human annotation by aligning data sample $x$ to $\mathcal{C}_{x}$ Human annotation by aligning data sample $x$ to $\mathcal{G}_{x}$
One-stage annotation: from $x$ to $\mathcal{C}_{x}$ Two-stage annotation: from $x$ to $\mathcal{G}_{x}$ and to $\mathcal{C}_{x}$
Opaque human annotation adherence to $\mathcal{G}$ Transparent human annotation adherence to $\mathcal{G}$
Coarse-grained evaluation based on $\mathcal{C}_{x}$ Fine-grained evaluation based on $\mathcal{G}_{x}$ and $\mathcal{C}_{x}$
Inconsistency between annotation process and model training Consistency between annotation process and model training
Allow for a single task formulation (one class set) Allow for multiple task formulations (multiple class sets)
Table 2: Comparison between the standard and the Guideline-Centered annotation processes.
Inconsistent Supervised Model Training.
In supervised learning, we train a machine learning (ML) model on annotated datasets to mimic the mapping from data samples $x$ to class subsets $\mathcal{C}_{x}$ performed by human annotators Hastie et al. (2009). However, this intent is not reflected in the training procedure nor in the model evaluation. During training, a data-driven classifier is trained to predict the class subset $\mathcal{C}_{x}$ corresponding to data sample $x$ without being provided with any knowledge of the guideline set $\mathcal{G}$. The underlying assumption is that the optimal mapping from $x$ to $\mathcal{C}_{x}$ learned by the model is the one adherent to the guideline set and thus similar to the human decision-making process. However, no explicit constraints prevent the model from learning spurious correlations from data and, more in general, exploiting shortcuts Geirhos et al. (2020). Moreover, since the guideline set is not part of the ML model workflow, there is no way to measure its adherence to the guideline set a-posteriori and, thus, we cannot validate the results under this lens.
3 Guideline-Centered Annotation
In Section 2, we discussed the shortcomings of the standard data annotation process, which, given a rich set of annotation guidelines $\mathcal{G}$ describing a task, produces only sample-specific class subset $\mathcal{C}_{x}$ with no explicit information on the underlying decision-making process grounded on $\mathcal{G}$. To address the presented issues, we propose a novel annotation process, denoted as Guideline-Centered (GC), aimed at enriching the annotation by keeping track of the guidelines associated with data samples. In this section, we introduce GC (Section 3.1) and describe its key properties (Section 3.2) by first recalling the standard annotation process under the lens of a novel formalization.
3.1 Formalization
We provide a formalization of the standard annotation process to set the ground for our GC annotation process.
Standard Annotation Process.
In the standard annotation process, annotators are provided with (i) the guideline set $\mathcal{G}$ describing the task, (ii) the class set $\mathcal{C}$ by which to annotate data, and (iii) the mapping existing between $\mathcal{G}$ and $\mathcal{C}$ linking each class $c$ with one or more guideline definitions $g$. Based on this information, annotators identify the class subset for each data sample. Formally, the annotator’s decision-making process can be seen as a function that aligns a data sample $x$ to its class subset $\mathcal{C}_{x}$:
$f:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{C})$ (1)
with $\mathcal{P}(\mathcal{C})$ the power set of the class set. Since the annotator is not instructed to report the guideline subset they rely on for annotating each data sample, the decision-making function $f$ is a black-box model proper of each annotator in which adherence to guidelines cannot be verified.
Guideline-Centered Annotation Process.
Conversely to standard annotation, in GC, annotators are provided with the guideline set $\mathcal{G}$ only and have no access to the class set $\mathcal{C}$ and the mapping between $\mathcal{G}$ and $\mathcal{C}$. GC annotation is composed of two stages: first, annotators are required to identify the guideline subset $\mathcal{G}_{x}\subseteq\mathcal{G}$ for each data sample $x$; then, the annotation process is finalized by mapping each $\mathcal{G}_{x}$ to the corresponding class set $\mathcal{C}_{x}$ based on a predefined mapping function. Formally, GC can be seen as the composition of two functions:
$h$ $:\mathcal{X}\rightarrow\mathcal{P}(\mathcal{G})$ (2)
$r$ $:\mathcal{G}\rightarrow\mathcal{C}$ (3)
[…]
3.2 Properties
In the following, we describe four major properties of GC addressing the issues of standard annotation highlighted in Section 2.1. Table 2 provides a comparison summary.
Transparent Annotation Adherence to Guidelines.
A key aspect of the GC annotation process is that annotators are unaware of the class set $\mathcal{C}$ as well as the class grounding function $r$. Given a data sample $x$, annotators focus solely on identifying and reporting the guideline subset $\mathcal{G}_{x}$ that relates with $x$. This means that unlike the standard annotation process where the focus is on the class set $\mathcal{C}$, GC is centered on the guideline set $\mathcal{G}$. This shift in focus is crucial. By not providing annotators with information about the class set or the class grounding function, the annotation process enforces strict adherence to the guideline set $\mathcal{G}$, which is essential in any standard annotation process for assessing its validity.
Fine-grained Annotation Evaluation.
As described in Section 2.1, discussion phases are essential for the annotation procedure, as they serve the purpose of reaching an agreement on edge cases. In GC, discussions revolve exclusively around the reported guideline subsets and do not involve any reasoning on the class set, as annotators do not have access to it during the annotation phase. In this setting, the convergence of annotators’ points of view can be effectively evaluated by inspecting the sample-specific guideline subsets they select. In addition, consider that the guideline set $\mathcal{G}$ contains far more information compared to the class set $\mathcal{C}$, as guidelines are textual definitions while the class set is a set of symbols. Thus, having access to the sample-specific guideline subset $\mathcal{G}_{x}$ allows annotation quality measures in terms of not only class subsets but also guideline subsets, unlocking a finer-grained analysis of annotated data.
Different Task Formulations.
GC is a two-stage annotation process composed of (i) a human annotation stage, where annotators apply their decision-making function $h$ to identify the guideline subset $\mathcal{G}_{x}$ for each data sample $x$, and (ii) a class grounding stage, in which the class subset $\mathcal{C}_{x}$ is retrieved from $\mathcal{G}_{x}$ by applying the class grounding function $r$. General examples of class grounding are class sets for multi-class and multi-label classification. The decomposition of the annotation process in these two phases enhance the decoupling of $\mathcal{G}$ and $\mathcal{C}$. As a result, we can use the annotated dataset created during the human annotation stage for different task formulations that share the same guidelines set. These formulations may stem from a taxonomy Kirk et al. (2023) or account for discrepancies in how a task should be addressed, as in the case of subjectivity detection described in Antici et al. (2024). In Figure 2, we report an example with three different class sets $\mathcal{C}^{\alpha},\mathcal{C}^{\beta},\mathcal{C}^{\gamma}$ corresponding to three possible formulations of the hate-speech detection task. The data sample $x$ is annotated with its guideline subset by a human annotator. The resulting $\mathcal{G}_{x}$ is mapped to the three class subsets by applying the class grounding function linking the set of guidelines $\mathcal{G}$ to the corresponding class set. As a result, we have three different class subsets $\mathcal{C}_{x}^{\alpha}$, $\mathcal{C}_{x}^{\beta}$, and $\mathcal{C}_{x}^{\gamma}$ at the cost of a single human annotation stage. This is a key advantage of the GC process over the standard annotation process, in which changing the class set would require fully re-annotate data.
Figure 2: GC annotation process allows defining different class sets ($\mathcal{C}^{\alpha}$, $\mathcal{C}^{\beta}$, $\mathcal{C}^{\gamma}$) and different class grounding functions ($r^{\alpha}$, $r^{\beta}$, $r^{\gamma}$) for the same guideline set $\mathcal{G}$ at the cost of a single human annotation stage (blue box). The annotator is unaware of the mapping between $\mathcal{G}$ and the class sets $\mathcal{C}^{\alpha}$, $\mathcal{C}^{\beta}$ and $\mathcal{C}^{\gamma}$.
Consistent Supervised Model Training.
As pointed out in Section 2.1, the guideline set is not part of a standard ML model training procedure, causing a discrepancy between human and automated decision-making processes. GC annotation process overcomes this limitation, allowing the training of a supervised ML model that predicts the sample-specific guideline subsets instead of the class subsets. This training setting is consistent with the human annotation stage, as the ML model focuses on predicting the guideline subsets $\mathcal{G}_{x}$ provided by the annotators. We highlight three main advantages of supervising a model with $\mathcal{G}_{x}$ compared to $\mathcal{C}_{x}$. First, we effectively prevent the model from learning spurious correlations between data samples and class subsets, as the model has no access to $\mathcal{C}_{x}$ during training. Second, the decoupling between $\mathcal{G}$ and $\mathcal{C}$ allows using the ML model for different task formulations, as its predictions can be mapped to different class sets relying on the corresponding class grounding functions. Third, we can rigorously evaluate the trained model’s performance by assessing its alignment with the defined guidelines. This involves analyzing disparities between predicted and ground-truth guideline subsets, fostering a deeper comprehension of the model’s decision-making mechanisms.
4 Related Work
Annotation paradigms have been recently proposed for data annotation to encode one shared view, such as the prescriptive paradigm Rottger et al. (2022), or different beliefs about a phenomenon, such as the descriptive Rottger et al. (2022), perspectivist Cabitza et al. (2023) paradigms. The ultimate goal of annotation paradigms is to regulate data annotation for a specific intended use of collected data. In the prescriptive paradigm, this objective is achieved by means of annotation guidelines Benikova et al. (2014); Sabou et al. (2014); Zaghouani et al. (2014); Krishna et al. (2023). Notable examples address subjective tasks, such as hate speech Salminen et al. (2019); Sap et al. (2022); Davani et al. (2023), stance detection Luo et al. (2020); AlDayel and Magdy (2021), sentiment analysis Waseem and Hovy (2016); Kenyon-Dean et al. (2018), and subjectivity detection Antici et al. (2024).
In all these approaches, annotators do not keep track of the annotation guidelines they rely on to annotate data samples. Therefore, although of crucial importance, the mapping between guidelines and data samples is lost during the process. To the best of our knowledge, Jikeli et al. (2023) is the only work that designs a data annotation process that enforces explicit grounding to annotation guidelines. In particular, the authors collect and annotate a corpus for the automated detection of antisemitic tweets. During the annotation process, annotators are required to specify which part of the definition of antisemitism applies to each data sample. Our work differs from the one of Jikeli et al. (2023) in several aspects. First, the work of Jikeli et al. (2023) is limited to a specific issue within social media content analysis and does not present a general formulation of the annotation process. Second, the proposed data annotation process relies on reporting both the guideline subset $\mathcal{G}_{x}$ and class subset $\mathcal{C}_{x}$ for each data sample $x$. This process suffers from an annotation shortcut where annotators may first define $\mathcal{C}_{x}$ and subsequently $\mathcal{G}_{x}$, thus reversing the intended annotation order. Furthermore, the annotation process is tight to a specific class set $\mathcal{C}$ and has increased annotation workload due to reporting both $\mathcal{G}_{x}$ and $\mathcal{C}_{x}$. We elaborate more on these two aspects in Section 5.
5 Discussion and Limitations
Alternative Annotation Methodology.
In Section 3.2, we discussed the benefits of GC methodology in reporting sample-specific guideline subset $\mathcal{G}_{x}$ compared to standard annotation procedure. As a possible alternative, we can consider an annotation process where the annotator reports both the guideline subset $\mathcal{G}_{x}$ and the class subset $\mathcal{C}_{x}$ for each data sample. This solution would allow for transparent adherence to annotation guidelines and fine-grained evaluation of annotations as GC methodology. However, this design choice is exposed to another form of annotation shortcut where the annotator may first determine $\mathcal{C}_{x}$ and then define $\mathcal{G}_{x}$ accordingly, thus reversing the expected annotation process. Furthermore, as the standard annotation process, this alternative is tailored to a specific class set $\mathcal{C}$, implying that a change in $\mathcal{C}$ requires re-annotating collected data. Lastly, reporting both $\mathcal{G}_{x}$ and $\mathcal{C}_{x}$ for a data sample $x$ carries the annotation cost of two distinct annotation processes, which is not desired.
Employing Fine-grained Class Sets.
Another point of discussion concerns the use of a fine-grained class set $\mathcal{C}$ instead of the guideline set $\mathcal{G}$. For instance, in Kirk et al. (2023), the authors explore different class sets for sexism detection, including one where each class $c$ has a one-to-one correspondence with a guideline definition $g$. While an annotation process based on such a rich class set $\mathcal{C}$ allows for a fine-grained annotation evaluation, it still suffers from two main weaknesses: annotation shortcut and support for single task formulation only. First, an annotator might still leverage their own belief to align a data sample $x$ to a class $c$ while bypassing its corresponding guideline definition $g$. As discussed in Section 2.1, the risk that an annotator views a class $c$ and its corresponding guideline definition $g$ as two distinct concepts during annotation, thus, leading to shortcuts where $g$ is ignored, cannot be verified since the annotation process is adherent to the guideline set $\mathcal{G}$ only implicitly. In contrast, in GC, the annotation process is slightly different, where strict adherence to the guideline set $\mathcal{G}$ is the pivotal point of the annotation. Thus, the risk of annotation shortcut is prevented. Second, while a fine-grained class set allows for several potential class groundings to more coarse-grained class sets (e.g., class sets organized in a taxonomy as in Kirk et al. (2023)), it still limited in its flexibility in addressing different task formulations. For instance, it is not possible to account for a change of the class grounding function $r$ while keeping the same class set $\mathcal{C}$ as in the subjectivity detection example discussed in Section 3.2 Antici et al. (2024).
Model Training with Guideline Definitions.
In recent years, several works have highlighted the importance of leveraging the textual description corresponding to a class $c$ in addressing a text classification problem, whether in the form of class names, short descriptions, or task instructions. These approaches differ in how class information is integrated and the particular learning method, whether through embedding-based enrichment Wang et al. (2018); Obeidat et al. (2019); Guo et al. (2021); Ozmen et al. (2023); Yin et al. (2023), metric learning guidance Wohlwend et al. (2019); Zhang et al. (2023); Nghiem et al. (2024), semantic retrieval Ruggeri et al. (2022, 2024), and by prompting LLMs He et al. (2023); Sainz et al. (2023); Tan et al. (2024), to name a few. While these approaches demonstrate that integrating class textual descriptions achieves superior classification performance, they still rely on the class set $\mathcal{C}$ for model training supervision. As discussed in Section 3.2, this formulation suffers from unconstrained model training, where the adherence of the data-driven training approach to the task conveyed by the guideline set $\mathcal{G}$ is nuanced. We argue that a shift towards training a model on predicting sample-specific guideline subsets $\mathcal{G}_{x}$ is needed to instruct and evaluate models in a way that is explicitly related to how a task is conceived, where an analysis through class symbols $c$ is always possible. We leave the evaluation of data-driven models on reporting $\mathcal{G}_{x}$ as future work.
Human Annotation Overhead.
In GC, there may be an initial overhead in time for human annotation with respect to the standard data annotation process. This is due to the requirement for annotators to report the guideline set, which is usually far more fine-grained than the class set and, thus, more time-consuming to internalize. To mitigate the annotation overhead, the guidelines should be designed for clarity and ease of adoption to assist annotators in effectively internalizing them and reducing the annotation overhead.
Other Annotation Paradigms.
The underlying idea of GC on keeping track of the guideline subset $\mathcal{G}_{x}$ associated with a data sample $x$ rather than its corresponding class subset $\mathcal{C}_{x}$ could, in principle, be extended to other annotation paradigms like the perspectivist one Cabitza et al. (2023). Here, several beliefs of distinct annotators could be grounded on guideline definitions $g$ rather than classes $c$, allowing for a more cohesive and fine-grained discussion and evaluation of collected annotations. More in general, the advantages of GC rely on the availability and definition of a guideline set depicting the rich spectrum of properties of a task. Therefore, its application to annotation paradigms where this requirement is not immediately satisfied, like in the descriptive one Rottger et al. (2022), is not straightforward. We believe that the applicability of GC to other annotation paradigms deserves a separate study and leave this analysis as future work.
Ethical Considerations.
An inherent problem of using guidelines in annotation processes is the potential bias they introduce. Annotators are often directed to follow prescribed criteria, which may lead to the inadvertent rejection of alternative, potentially valid annotation criteria. The prescriptive paradigm assumes that the guidelines comprehensively cover all possible cases, which is rarely the case in practice. Literature work highlights the necessity for iterative processes in annotation, emphasizing the need for continual refinement and adaptation of guidelines (e.g., Lee et al. (2023); Guzmán-Monteza (2023)). This iterative approach allows for the inclusion of diverse perspectives and mitigates the ethical concern of bias, promoting a more robust annotation framework.
6 Conclusions
In this paper, we presented the Guideline-Centered methodology, a novel annotation process focused on reporting the annotation guidelines associated with each data sample. We described the advantages of our methodology in comparison with the standard data annotation process performed according to the prescriptive paradigm. These advantages include unlocking a finer-grained analysis of annotated data, allowing multiple task formulations at the cost of a single human annotation phase, and consistency between data annotation and model training where both processes are grounded on annotation guidelines. We think that our methodology can foster researchers in developing high-quality benchmarks for discerning the decision-making process and the limitations of state-of-the-art models, an ever-recurring need in NLP Yadkori et al. (2024).
As future work, we identify two main research directions. First, considering the hot research topic of large language models (LLMs) evaluation Zhao et al. (2023); Minaee et al. (2024), we plan to evaluate LLMs in reporting guideline definition subsets ($\mathcal{G}_{x}$) for several tasks. Second, we aim to extend our analysis to other annotation paradigms like the descriptive one Rottger et al. (2022).